# Adapting Large Language Models for Character-based Augmentative and Alternative Communication

### v2

@@ Abstract
Users of Augmentative and Alternative Communication (AAC) may write letter-by-letter via an interface that uses a character language model. However, most state-of-the-art large pretrained language models predict subword tokens of variable length. We investigate how to practically use such models to make accurate and efficient character predictions. We fine-tune models using a large dataset of sentences we curated in which each sentence is rated according to how useful it might be for spoken or written AAC communication. We find that using an algorithm to produce character predictions from a subword large language model provides more accurate predictions than adding a classification layer or using a byte-level model. We also find that our domain adaptation procedure is effective at improving model performance on simple, conversational text.
@@
Dylan Gaines∗ Michigan Technological University Houghton, MI, USA dcgaines@mtu.edu Keith Vertanen∗ Michigan Technological University Houghton, MI, USA vertanen@mtu.edu
@@ 1 Introduction
Augmentative and Alternative Communication (AAC) devices allow non-speaking individuals to communicate in face-to-face conversations and asynchronously via messaging systems such as email. Depending on their disability, AAC users may be unable to use a conventional computer keyboard and may instead rely on alternative input methods such as eye tracking or triggering a single switch (e.g. by twitching a muscle or puffing air). Unfortunately, such input methods can make writing slow and tiring. The writing speed of non-speaking AAC users varies between studies and methods, but is frequently less than 10 words per minute (WPM), with an average closer to 5 WPM Polacek et al. (2017).
While some AAC users may engage in long conversations, communications are often shorter and more isolated in nature due to input rate limitations. AAC users with severe motor impairments may write letter-by-letter using an interface such as Dasher Ward and MacKay (2002), the RSVP Keyboard Orhan et al. (2012), or Nomon Broderick and MacKay (2009). These AAC interfaces all require the probability over the next character based on a user’s previous text.
We wanted to know if recent advances in pretrained large language models (LLMs) really help with the “simple” case of predicting the next letter in a sentence with no prior context. Numerous papers have shown that ever-larger models such as Llama 3 Grattafiori et al. (2024), Mistral 7B Jiang et al. (2023), and GPT-4 OpenAI et al. (2024)) provide gains on more “complex” natural language processing (NLP) tasks such as question answering, sentiment analysis, and machine translation.
@@
We have made our algorithm’s source code, classified data sets, our classification model, and our best language models available.111https://osf.io/ajm7t/
This paper is structured as follows: first, we motivate why character predictions are important for AAC text entry (Section 2) and describe our algorithm for making such predictions using a subword tokenized LLM (Section 3). Next, we find the best starting model, testing performance on a private collection of AAC-like text (Section 4). We then explore adapting the LLM using small in-domain sets and using text classified from much larger out-of-domain sets (Section 5). Next, we compare our search algorithm against other possible approaches (Section 6). Finally, we show our results on five test sets (Section 7) before wrapping up with discussion (Section 8) and conclusions (Section 9).
@@ 2 Related work
2.1 Language modeling for AAC devices
Language models have been frequently studied in AAC applications in an attempt to improve communication rates through, for example, word completions and predictions. Word completions attempt to predict a user’s intended current word before they finish typing it. Word predictions attempt to predict subsequent words. Kristensson and Müllners (2021) show that word predictions can increase or decrease a user’s entry rate depending on how and when predictions are displayed and the user’s typing strategy. Yusufali et al. (2023, 2024) used LLMs to generate word predictions for AAC devices and showed entry rates up to 30.4 WPM.
Retrieval techniques allow users to store sentences or phrases that they might frequently use Todman et al. (2008). These messages can then be retrieved by the user by entering specific keywords Langer and Hickey (1998) or matching context tags Kristensson et al. (2020). The drawback to retrieval is that the utterances must be pre-stored by the user in their device, and they are not as adaptable to different conversational situations.
Recent advances in LLMs have allowed researchers to develop more dynamic abbreviation expansion interfaces. Cai et al. (2022, 2023) developed a system where users abbreviate each word to its first letter, and it leverages the context of the utterance to decode the intended expansion with an LLM. KWickChat Shen et al. (2022) allows the user to enter a few keywords and uses an LLM to generate expanded utterances based on the provided keywords, some contextual details about the user, and the dialogue from their conversational partner. Valencia et al. (2023) also developed an abbreviation expansion system that expands single words into complete sentences using conversational context and background information about the user.
@@ 2.2 Applications for character predictions
While these approaches can help accelerate input with some AAC devices, other systems such as brain-computer interfaces (BCIs) rely on noisier signals. P300 spellers Farwell and Donchin (1988) present visual stimuli to users and measure their response using EEG electrodes. One example is the RSVP Keyboard Orhan et al. (2012), which shows a rapid series of characters to try and determine a user’s intended character. Using RSVP, it can take several series worth of user responses to gain enough evidence to type a character. RSVP uses a character n-gram language model in combination with the EEG evidence to reduce the number of series needed to select a character. We hypothesize a LLM will provide more accurate probabilities and reduce the number of series needed.
Nomon Broderick and MacKay (2009); Bonaker et al. (2022b, a) allows users to select words or characters by activating a switch when the rotating clock corresponding to their target reaches noon. It can take several clicks to make a single selection since the system needs to differentiate between similar clocks. Since Nomon provides both word and character selection options, it needs a language model that can evaluate the likelihood of each possibility to help accelerate writing.
Dasher Ward and MacKay (2002) is another interface that relies on character-level language model predictions. In Dasher, the letter selection options are displayed vertically on the right side of the screen. To enter text, users move a cursor through the sequence of nested letter boxes corresponding to their desired text. Dasher can be controlled by various devices such an eye-tracker, mouse, or touchscreen. As users progress through their text, more likely letter options are displayed as larger targets, making it easier to select them.
Dasher uses the prediction by partial match (PPM) algorithm to determine letter probabilities. PPM is similar to an n-gram language model but has a data structure making it easy for to continually adapt to a user’s writing. We hope to improve on this using large language models. However, a key challenge would be generating predictions quickly enough to support Dasher’s visualization of the many possible future character sequences.
@@ 3 Character prediction algorithm
As we discussed previously, many modern large transformer models use subword tokens instead of single characters or full words. The first step in using a subword token language model is to convert the context into a sequence of tokens via tokenization. Each model comes with its own tokenizer, which uses a greedy algorithm (e.g. byte pair encoding Sennrich et al. (2016)) to segment any given text into tokens in its model’s vocabulary.
This subword tokenization complicates the process of obtaining a distribution of character probabilities since adding another character to the end of a section of text could (1) begin a new token containing only the new character, (2) extend the existing most recent token with the new character, or (3) rearrange the optimal tokenization of the most recent word. For the last case, take for example writing the word “yesterday”. If the user has written “yester” thus far, the most likely tokenization may be into “y-este” (where - denotes token boundaries). Adding an “r” could result in the most likely tokenization changing to “yes-ter”.
To properly consider each of these possibilities, we can remove one or more tokens from the end of the context when making predictions to allow the model the option of extending it, as opposed to only adding a new token. Removing more tokens allows the model more flexibility to rearrange the optimal tokenization and may lead to better predictions, but at the cost of a higher prediction time. In this work, we remove all tokens in the current word to allow maximal flexibility in tokenizing the partial word. Going further back would provide no benefit because, in all models we used, the space character only appears at the start of tokens.
We then ask the model to estimate the likelihood of the next token and evaluate any token that matches our context. For efficiency, we only track a certain number of hypotheses at a time, known as the beam width. We continue expanding each hypothesis until it has regenerated the removed context tokens plus at least one character. We then store the likelihood for each final prediction in a list based on the character that directly follows the context. We continue this process until there are no more hypotheses to extend, or, for efficiency, the number of completed hypotheses exceeds a max completed parameter. Then we sum the likelihoods stored for each character in our symbol set and normalize to sum to one, arriving at a distribution over our symbol set. See Appendix C for our pseudocode search algorithm or our supplementary materials for a Python implementation.
@@ 4.1 Private AAC-like test data
Unfortunately, actual messages from AAC users are difficult to source due to practical, ethical, and privacy reasons. There are a number of datasets that have been used to approximate AAC communications such as voice banking messages Costello (2014a), Switchboard Godfrey et al. (1992), DailyDialog Li et al. (2017), COMM2 Vertanen (2013), and Turk dialogues Vertanen (2017). However, all of these datasets are available online and may have been used in training LLMs.
As we wish to fairly compare performance between different LLMs, as well as between LLMs and an n-gram language model optimized for AAC-like text, we created two private tests sets, one for informal written communications (i.e. similar to mobile messaging), and one for person-to-person communications222We intend to provide access to researchers on the condition they not leak the text online.. We denoted the two distinct genres as written and spoken respectively. We sourced our text from compositions made by workers on Amazon Mechanical Turk in response to various prompts. For details of how we constructed our test sets, including example sentences, see Appendix A.
We split each genre in half, creating a written dev set (8.6 K words), a written test set (9.0 K words), a spoken dev set (11.3 K words), and a spoken test set (10.8 K words). We will use the dev sets for most of our experiments, reserving the test sets for a final evaluation (Section 7). Our final evaluation will also use text written by an AAC user, messages written by people with ALS for voice banking Costello (2014b), and responses to common conversational situations Vertanen (2013).
@@ 4.2 Perplexity experiments
We converted our private test sets to lowercase and removed anything not a letter, apostrophe, or space. We calculated per-character perplexity (including spaces) on our dev sets without any sentence end token. Sentences were conditioned on the sentence start token appropriate for the model under test.
We used 16-bit floating point inference for the LLMs. We used our search algorithm (Section 3) for LLM predictions. We found a search beam of 8 and a max completed stopping criteria of 32 K provided a perplexity close to the perplexity without pruning while significantly speeding inference.
We compare against a baseline of a 12-gram language model optimized for AAC-like text, denoted AAC n-gram.333https://osf.io/c6mnz This model was trained using Witten-Bell smoothing on 21 B characters from Common Crawl, Reddit, movie subtitles, and Twitter. Training sentences were selected via cross-entropy difference selection Moore and Lewis (2010) using in-domain models trained on DailyDialog Li et al. (2017), short emails, and AAC-like messages Vertanen and Kristensson (2011).
As shown in Table 1, per-character perplexity decreased with increasing model size, though gains plateaued for the OPT family Zhang et al. (2022) of models around 1.3 B parameters. Thus, even given the “easy” nature of our NLP task (i.e. predicting the next character within an isolated sentence), bigger models were often better. While the n-gram model optimized for AAC did better than DistilGPT with a similar number of parameters, larger LLMs eventually outperformed it despite the LLMs not being fine-tuned on AAC-like text.
@@
Determining the cause of the perplexity differences between model families is difficult due to the many differences in the underlying LLM architectures, training data, and training procedures. But given that the performance of n-gram models on AAC-like text have been shown to depend on the training data Vertanen and Kristensson (2011), we conjecture difference may be due to the amount of well-matched English text versus other types of text (e.g. other natural or programming languages).
Our n-gram model inferences using KenLM Heafield (2011) were fast at 0.1 ms per prediction (using a CPU). Despite substantial performance optimization efforts as well as using an A100 GPU for inference, our search algorithm was much slower at 38–93 ms per prediction. In general, larger models took longer. We will investigate other models for speeding inference in Section 6.
@@
| distilgpt2 | 82 | 2.71 | 37.5 |
| gpt2 | 124 | 2.54 | 42.2 |
| gpt2-medium | 355 | 2.45 | 52.6 |
| gpt2-xl | 1558 | 2.38 | 74.2 |
| opt-125m | 125 | 2.39 | 46.6 |
| opt-350m | 331 | 2.32 | 61.1 |
| opt-1.3b | 1316 | 2.22 | 60.9 |
| opt-2.7b | 2652 | 2.21 | 70.3 |
| opt-6.7b | 6658 | 2.20 | 70.5 |
| opt-13b | 12853 | 2.23 | 91.7 |
| Meta-Llama-3.1-8B | 8030 | 2.31 | 93.4 |
@@ 5 Optimizing language models for AAC
Our results in Table 1 were on lowercase versions of our test sentences. Our search algorithm sums over all matching following tokens ignoring case. However, we conditioned predictions on all lowercase context. This simulates an AAC interface where the user does not bother to denote case, or when casing is added later in some way. We found we could slightly improve perplexity on our dev sets using opt-350m from 2.32 to 2.30 by upper casing the first letter of the sentence and the words: I, I’m, I’ll, I’ve, and I’d. We found this simple approach performed similar to using the human-supplied case. We will use this automatic casing method for the remainder of our experiments.

### v3

@@ Abstract
Users of Augmentative and Alternative Communication (AAC) may write letter-by-letter via an interface that uses a character language model. However, most state-of-the-art large pretrained language models predict subword tokens of variable length. We investigate how to practically use such models to make accurate and efficient character predictions. Our algorithm for producing character predictions from a subword large language model (LLM) provides more accurate predictions than using a classification layer, a byte-level LLM, or an n-gram model. Additionally, we investigate a domain adaptation procedure based on a large dataset of sentences we curated based on scoring how useful each sentence might be for spoken or written AAC communication. We find our procedure further improves model performance on simple, conversational text.
@@ 1 Introduction
Augmentative and Alternative Communication (AAC) devices allow non-speaking individuals to communicate in face-to-face conversations and asynchronously via messaging systems such as email. Depending on their disability, AAC users may be unable to use a conventional computer keyboard and may instead rely on alternative input methods such as eye tracking or triggering a single switch (e.g. by twitching a muscle or puffing air). Unfortunately, such input methods can make writing slow and tiring. The writing speed of non-speaking AAC users varies, but is frequently less than 10 words per minute (WPM), with an average closer to 5 WPM Polacek et al. (2017).
While some AAC users may engage in long conversations, communications are often shorter and more isolated in nature due to the input rate limitation. AAC users with severe motor impairments may write letter-by-letter using an interface such as Dasher Ward and MacKay (2002), the RSVP Keyboard Orhan et al. (2012), or Nomon Broderick and MacKay (2009). These AAC interfaces all require the probability distribution over the next character based on a user’s previous text.
We wanted to know if recent advances in pretrained large language models (LLMs) really help with the “simple” case of predicting the next letter in a sentence with only that sentence as context. Numerous papers have shown that ever-larger models such as Llama 3 Grattafiori et al. (2024), Mistral 7B Jiang et al. (2023), and GPT-4 OpenAI et al. (2024) provide gains on more “complex” natural language processing tasks such as question answering, sentiment analysis, and machine translation.
@@
We have released our algorithm’s source code, our classified data sets, our classification model, and our best language models as part of this work.111https://osf.io/ajm7t/
This paper is structured as follows: first, we motivate why character predictions are important for AAC text entry (Section˜2) and describe our algorithm for making such predictions using a subword tokenized LLM (Section˜3). Next, we find the best starting model, testing performance on a private collection of AAC-like text (Section˜4). We then explore adapting the LLM using small in-domain sets and using text classified from much larger out-of-domain sets (Section˜5). Next, we compare our search algorithm against other possible approaches (Section˜6). Finally, we show our results on five test sets (Section˜7) before wrapping up with discussion (Section˜8) and conclusions (Section˜9).
@@ 2 Related work
2.1 Language modeling for AAC
Language models have often been studied in AAC in hopes of improving communication speed. Work has focused on four main topics: word completions and predictions, sentence and phrase retrieval, abbreviation expansion, and domain adaptation.
Word completions and predictions. Word completions attempt to predict a user’s intended current word before they finish typing it, while word predictions attempt to predict subsequent words based on the likelihood of each word in the context of the user’s sentence. Prior work by Kristensson and Müllners (2021) showed that word predictions can increase or decrease a user’s entry rate depending on how and when predictions are displayed and the user’s typing strategy. Yusufali et al. (2023, 2024) used LLMs to generate word predictions for AAC devices and showed entry rates up to 30.4 WPM.
Sentence and phrase retrieval. Retrieval techniques allow users to store sentences or phrases that they frequently use Todman et al. (2008). These messages can then be retrieved by the user by entering specific keywords Langer and Hickey (1998) or matching context tags Kristensson et al. (2020). The drawback to retrieval is utterances must be pre-stored by the user in their device, so they are not as adaptable to different conversational situations.
Abbreviation expansion. Recent advances in LLMs have allowed the development of more dynamic abbreviation expansion interfaces. Cai et al. (2022, 2023) developed a system where users abbreviate each word to its first letter. The system leverages the context of the utterance and an LLM to decode the intended text. KWickChat Shen et al. (2022) allows the user to enter a few keywords and uses an LLM to generate expanded utterances based on the provided keywords, some contextual details about the user, and the dialogue from their conversational partner. Valencia et al. (2023) also developed an abbreviation expansion system that expands single words into complete sentences using conversational context and background information about the user.
Domain adaptation. The type of text an AAC user may want to write can differ from typical sources of language model training data (e.g. text from web pages). Vertanen and Kristensson (2011) used a collection of crowdsourced fictional AAC messages to filter sentences from large social media datasets using cross-entropy difference Moore and Lewis (2010). Their optimized n-gram language models improved perplexity and potential keystroke savings on AAC-like test sets. Adhikary et al. (2021) used the same crowdsourced collection to select sentences from text scraped from the web. They found using a BERT-based classifier to select sentences slightly outperformed selection via cross-entropy difference.
@@ 2.2 Applications for character predictions
While word and phrase-based language modeling approaches can help accelerate input with many AAC devices, some devices rely primarily on character-at-a-time input. In these systems, it may take several user interactions to input a single character due to a user’s noisy input signal. Character predictions from a language model can help accelerate input of individual characters in such systems.
Brain-computer interfaces. P300 spellers Farwell and Donchin (1988) present visual stimuli to users and measure their response using EEG electrodes. For example, the RSVP Keyboard Orhan et al. (2012) displays a rapid series of characters to users. Using RSVP, it can take several series worth of EEG responses to infer a user’s desired character. RSVP uses a character n-gram language model in combination with the EEG evidence to reduce the number of series needed. We hypothesize an LLM will provide more accurate probabilities and reduce the number of series needed.
Nomon. Nomon Broderick and MacKay (2009); Bonaker et al. (2022b, a) allows users to select words or characters by activating a switch when the rotating clock corresponding to their target reaches noon. It can take several clicks to make a single selection since the system needs to differentiate between similar clocks. Since Nomon provides both word and character selection options, it needs a language model that can evaluate the likelihood of each possibility to help accelerate writing.
Dasher. Dasher Ward and MacKay (2002) inherently relies on character-level language model predictions. In Dasher, the letters of the alphabet are displayed vertically on the right side of the interface. To enter text, users move their cursor through the sequence of nested letter boxes corresponding to their desired text. Dasher can be controlled by various devices such an eye-tracker, mouse, or touchscreen. As users progress through their text, more likely letter options are displayed as larger targets, making it easier to select them.
Dasher uses the prediction by partial match (PPM) algorithm Cleary and Witten (1984). PPM is similar to an n-gram model but has a data structure making it easy to continually adapt to a user’s writing. We hope to improve on this using LLMs. However, a key challenge is generating predictions quickly enough to support Dasher’s visualization of the many possible future character sequences.
@@ 3 Character prediction algorithm
As we discussed previously, many modern large transformer models use subword tokens instead of single characters or full words. The first step in using a subword token language model is to convert the context into a sequence of tokens via tokenization. Each model comes with its own tokenizer, which uses a greedy algorithm, such as byte pair encoding Sennrich et al. (2016), to segment any given text into tokens in its model’s vocabulary.
This subword tokenization complicates the process of obtaining a distribution of character probabilities since adding another character to the end of a section of text could (1) begin a new token containing only the new character, (2) extend the existing most recent token with the new character, or (3) rearrange the optimal tokenization of the most recent word. For the last case, take for example writing the word “yesterday”. If the user has written “yeste” thus far, the most likely tokenization may be into “y-este” (where - denotes token boundaries). Adding an “r” could result in the most likely tokenization changing to “yes-ter”.
To properly consider each of these possibilities, we must remove one or more tokens from the end of the context to allow the model the option of extending it, as opposed to only adding a new token. Removing more tokens allows the model more flexibility to rearrange the optimal tokenization and in our experience leads to better predictions, but at the cost of a higher prediction time. In this work, we remove all tokens in the current in-progress word to allow maximal flexibility in tokenizing the partial word. Going further back would provide no benefit because, in all models we used, the space character only appears at the start of tokens and thus tokens cannot span word boundaries.
We then ask the model to estimate the likelihood of the next token and evaluate any token that matches our context. For efficiency, we only track a certain number of hypotheses at a time, known as the beam width. We continue expanding each hypothesis until it has regenerated the removed context tokens plus at least one character. We then store the likelihood for each final prediction in a list based on the character that directly follows the context. We continue this process until there are no more hypotheses to extend, or, for efficiency, the number of completed hypotheses exceeds a max completed parameter. Then we sum the likelihoods stored for each character in our symbol set and normalize to sum to one, arriving at a distribution over our symbol set. See Appendix˜C for our pseudocode search algorithm or our supplementary materials for a Python implementation.
To give a concrete example of the benefit of removing tokens from the end of the context, we further examine the “yeste” example. If we do not remove any context, and simply marginalize the probabilities over the first character in each token, the probability of the next letter being R is 0.0047. Using our search algorithm with context removal, this jumps to 0.9999. We suspect this is due to the training process of the model; during training it doesn’t frequently encounter fragmented words and is allowed to tokenize the full word in the most optimal way. By not removing any context, we force the model to infer text with suboptimal tokenizations, causing it to struggle.
@@ 4.1 Private AAC-like test data
Unfortunately, actual text from AAC users is difficult to source due to practical, ethical, and privacy reasons. There are a number of datasets that have been used to approximate AAC communications such as voice banking messages Costello (2014), Switchboard Godfrey et al. (1992), DailyDialog Li et al. (2017), COMM2 Vertanen (2013), and Turk dialogues Vertanen (2017). However, these datasets are available online and may have been used in training the LLMs we want to evaluate.
As we wish to fairly compare performance between different LLMs, as well as between LLMs and an n-gram language model optimized for AAC-like text, we created two private tests sets, one for informal written communications (i.e. similar to mobile messaging), and one for person-to-person communications. We denoted the two distinct genres as written and spoken respectively.222We intend to provide access to researchers on the condition they not leak the text online or to hosted models. We sourced our text from compositions made by workers on Amazon Mechanical Turk in response to various prompts. For details of how we constructed our test sets, including example sentences, see Appendix˜A.
We split each genre in half, creating a written dev set (8.6 K words), a written test set (9.0 K words), a spoken dev set (11.3 K words), and a spoken test set (10.8 K words). We use the dev sets for most of our experiments, reserving the test sets for a final evaluation (Section˜7). Our final evaluation also uses text written by an AAC user, messages written by people with ALS for voice banking Costello (2014), and responses to common conversational situations Vertanen (2013).
@@ 4.2 Perplexity experiments
We converted our private test sets to lowercase and removed anything not an English letter, apostrophe, or space. We calculated per-character perplexity (including spaces) on our dev sets without any sentence end token and using the sentence start token appropriate for the model under test.
We used 16-bit floating point inference for the LLMs. We used our search algorithm (Section˜3) for LLM predictions. We found a search beam of 8 and a max completed stopping criteria of 32 K provided a perplexity close to the perplexity without pruning while significantly speeding inference.
We compared against a baseline of a 12-gram language model optimized for AAC-like text, denoted AAC n-gram.333https://osf.io/c6mnz This model was trained using Witten-Bell smoothing on 21 B characters from Common Crawl, Reddit, movie subtitles, and Twitter. Training sentences were selected via cross-entropy difference selection Moore and Lewis (2010) using in-domain models trained on DailyDialog Li et al. (2017), short emails, and AAC-like messages Vertanen and Kristensson (2011).
As shown in Table˜1, per-character perplexity decreased with increasing model size, though gains plateaued for the OPT model family Zhang et al. (2022) at around 1.3 B parameters. Thus, even given the “easy” nature of our task (i.e. predicting the next character within an isolated sentence), bigger models were often better. While the n-gram model optimized for AAC did better than DistilGPT with a similar number of parameters, larger LLMs eventually outperformed it despite the LLMs not being fine-tuned on AAC-like text.
@@
Determining the cause of the perplexity differences between model families is difficult due to the many differences in the underlying LLM architectures, training data, and training procedures. But given that the performance of n-gram models on AAC-like text have been shown to depend on the training data Vertanen and Kristensson (2011), we conjecture the difference may be due to the amount of well-matched English text versus other types of text (e.g. other natural or programming languages).
Our n-gram model inferences using KenLM Heafield (2011) were fast at 0.1 ms per prediction (using a CPU). Despite substantial performance optimization efforts as well as using an NVIDIA A100 GPU for inference, our search algorithm was much slower at 55–120 ms per prediction. In general, larger models took longer. We will investigate other methods for faster inference in Section˜6.
@@
| distilgpt2 | 82 | 2.71 | 55.0 |
| gpt2 | 124 | 2.54 | 60.0 |
| gpt2-medium | 355 | 2.45 | 70.1 |
| gpt2-xl | 1558 | 2.38 | 91.7 |
| opt-125m | 125 | 2.39 | 64.3 |
| opt-350m | 331 | 2.31 | 78.6 |
| opt-1.3b | 1316 | 2.22 | 77.9 |
| opt-2.7b | 2652 | 2.21 | 86.8 |
| opt-6.7b | 6658 | 2.20 | 89.5 |
| opt-13b | 12853 | 2.23 | 108.8 |
| Meta-Llama-3.1-8B | 8030 | 2.31 | 119.9 |
| Meta-Llama-3.2-1B | 1236 | 2.37 | 97.0 |
| Meta-Llama-3.2-3B | 3213 | 2.33 | 115.0 |
@@ 5 Optimizing language models for AAC
Our results in Table˜1 were on lowercase versions of our test sentences. Our search algorithm sums over all matching following tokens ignoring case. However, we conditioned predictions on all lowercase context. This simulates an AAC interface where the user does not bother to denote case, or when casing is added later in some way. We found we could slightly improve perplexity on our dev sets using opt-350m from 2.31 to 2.30 by upper casing the first letter of the sentence and the words: I, I’m, I’ll, I’ve, and I’d. We found this simple approach performed similar to using the human-supplied case. We use this automatic casing method for the remainder of our experiments.